OpenStudy (anonymous):

write an equation for the are A in terms of L with a perimeter of 60

OpenStudy (anonymous):

is this a square

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im usually good at this stuff but my new teacher doesnt explain things very well

OpenStudy (anonymous):

well then area of what?

OpenStudy (anonymous):

your supposed to write an equation for the are of any rectangle with the area of 60 meters

OpenStudy (anonymous):

area* perimeter*

OpenStudy (anonymous):

perimeter of 60 meters

OpenStudy (anonymous):

please help this is what i have so far 60-2L=2w so w=30-w

OpenStudy (anonymous):

so you have w=30-l A=l*w substitute (30-l) for w so now we have A=l*(30-l) which simplified is A=30l-l^2